# Exploring LLM-based Student Simulation for Metacognitive Cultivation

### v2

@@ 4.3 Agent Profile Distribution
Among the Big Five personality traits, agents with low extraversion and high agreeableness are more prevalent in the final selection. It is because high extraversion often indicates strong energy and positivity, which may not align well with typical characteristics of students experiencing learning difficulties. In contrast, high agreeableness is more frequently observed, as agents tend to exhibit affirmative responses, making agents with high agreeableness more likely to receive higher alignment scores. For the study questionnaire, q2 (Do you have trouble taking tests or completing assignments?) shows a significantly higher proportion of "Yes" responses in the selected agents. It suggests that LLM-based student simulations better capture students with such difficulties, aligning with real-world academic advising practices where these challenges are common. Notably, in the four trait questionnaire, the proportion of agents with strong self-efficacy and self-regulation increases substantially after filtering. It is because agents inherently display high self-efficacy and structured planning abilities in conversations, even though these traits may not typically align with students facing learning difficulties. Nevertheless, such agents still achieve high scores in behavior consistency evaluations.
@@ 4.4 Influence of Features in Profile
To investigate the influence of the features in the agent profile on the final profile and behavior scores, we performed a hot encoding of the features of the profile and used a Random Forest model (Breiman, 2001) to fit the propagated profile score and behavior score, where the size of test sets is 20% and the number of trees is 100. We then ranked the features based on their relative importance(Archer and Kimes, 2008), where the relative importance of each feature is calculated by dividing its importance score by the highest among all.
@@ 4.5 Case Study
We showcase LLM scoring examples in Figure 7, comparing a good case and a bad case in profile and behavior evaluation, accompanied by expert scores, agreement levels, and explanations. In the good case (top), the questioning agent $Q$ detects an inconsistency in the student’s profile—17 years old while enrolled as a third-year Master’s student. The student agent fails to provide a compelling justification for this contradiction. Additionally, the LLM accurately assesses the student agent’s self-reported academic pressure, where the reasoning behind the discrepancy between daily and long-term pressure is somewhat reasonable. Regarding behavior evaluation, the scorer correctly identifies that the student agent violates the role of being a student who need academic advising. It demonstrates the model’s sensitivity to instruction following and reveals its occasional struggle with adherence to long instructions. In contrast, the bad case (bottom) illustrates a scenario where the LLM’s evaluation deviates significantly from human judgment. The agent’s profile presents conflicting traits, such as experiencing stress while reporting no difficulty in tests and assignments. The LLM questions these inconsistencies but assigns a relatively low score due to its strict conflict detection. However, expert annotations suggest that such contradictions are not inherently erroneous, but reflect common real-world cases where students manage stress effectively despite academic challenges. Furthermore, the LLM scorer underperforms behavior evaluation by misinterpreting the agent’s conscientiousness and self-efficacy, leading to lower agreement scores with human experts. The results indicate that while LLM scorers is adept at identifying explicit conflicts, it may over-penalize cases where nuanced human explanations exist.
@@ 5 Conclusion
Our work explores the potential of automatically generating and filtering high-quality simulated student agents to enhance the authenticity of learning difficulty simulations. However, researchers and practitioners must consider several limitations and risks. First, our approach focuses on prompt-based student simulation, particularly extending the scope to metacognition. We do not explore more complex agent simulation frameworks like workflow-based or retrieval-augmented generation (RAG) simulations. These approaches will be explored in future work to improve simulation fidelity further. Second, some studies evaluate LLM-based human simulation by Turing tests (Aher et al., 2023), measuring the model’s success rate in mimicking human-like responses. Such evaluations emphasize linguistic style. Our work prioritizes metacognitive abilities and learning limitations in simulation rather than stylistic authenticity. We explicitly addressed this distinction during expert annotation to ensure the evaluation remains aligned with our research objectives. Third, our approach assumes that structured profiles can effectively capture student attributes and learning difficulties. However, real-world learning behaviors are highly dynamic and context-dependent. Future work should explore adaptive modeling techniques that better reflect the evolving nature of student cognition and engagement.
@@
Additionally, ethical risks must be considered. The simulated student agents are generated based on training data, which may introduce biases, stereotypes, or unintended inaccuracies in student representations. These biases could affect the fairness and applicability of educational interventions. Researchers must remain vigilant in mitigating such risks through continuous evaluation, bias detection, and responsible deployment of simulated agents.
Appendix A Related Work
A.1 LLM-based Human Simulation
Recent studies on human simulation have increasingly leveraged large language models (LLMs) to emulate complex behaviors across social, economic, and physical domains. In the social realm, Gao et al. (2023) and Qian et al. (2024) demonstrated how LLM agents replicate interpersonal interactions and collaboration, enabling realistic group dynamics. From an economic perspective, Horton (2023) and Zhao et al. (2023) showed that LLM-based simulations can capture market behaviors and decision-making patterns aligned with established economic theories. In the physical domain, Jin et al. (2023) and Zou et al. (2023) employed LLM agents to model mobility and driving behaviors in virtual environments, highlighting human-like performance and adaptability. Hybrid approaches further integrate these dimensions, as illustrated by Park et al. (2023) and Williams et al. (2023), showcasing the potential of LLM-driven simulations to generate intricate, multifaceted human-like behavior. In the medical context, Du et al. (2024) introduced EvoPatient, an agent coevolution framework that models simulated patients and doctor agents engaging in multi-turn dialogues. Similar to simulating patients, academic advising, where an institutional representative provides guidance to a college student on academic, social, or personal matters (Kuhn, 2008), helping students maximize their college experience (Miller, 2011). However, to the best of our knowledge, little work explores simulating students in academic advising using LLM.
A.2 Student simulation in Education
The emergence of LLMs has spurred extensive research on simulat at the individual level(Yu et al., 2024; Zhang et al., 2024b). Markel et al. (2023) uses LLMs to simulate students to train teaching assistants (TAs) for office hours. It categorizes TA-student interactions, performs A/B testing to evaluate system effectiveness, and addresses academic and personal issues like study methods and stress management. Lu and Wang (2024) models student profiles by simulating their responses to multiple-choice questions, predicting their performance, and evaluating their conceptual understanding. It highlights how LLM-based simulations can assist teachers in assessing the quality of their test items. Liu et al. (2024) focuses on creating virtual students with diverse personality types (e.g., the Big Five or MBTI) to train teachers’ pedagogical skills. Jin et al. (2024) evaluates teaching agents by simulating students with varying traits and knowledge levels, then generating conversations to assess agent responses. Its features resemble debugging tools, enabling real-time agent performance analysis across student profiles. Ma et al. (2024) constructs conversational datasets based on real teacher-student interactions, fine-tuning LLMs to simulate diverse student types in language learning. The virtual students underwent a Turing test, where some personalities closely resembled real humans, benefiting from high-quality, real-student data for fine-tuning. Despite the success they have achieved, simulating students with discrepancies in knowledge or learning skills remains unexplored.
@@ Appendix B Experimental Settings
B.1 Settings
B.1.1 Configurations
B.1.2 Annotation Settings
B.2 Pipeline Performance Metrics
@@
[…]

### v3

@@ 4.4 Agent Profile Distribution
Among the Big Five personality traits, agents with low extraversion and high agreeableness are more prevalent in the final selection. It is because high extraversion often indicates strong energy and positivity, which may not align well with typical characteristics of students experiencing learning difficulties. In contrast, high agreeableness is more frequently observed, as agents tend to exhibit affirmative responses, making agents with high agreeableness more likely to receive higher alignment scores. For the study questionnaire, q2 (Do you have trouble taking tests or completing assignments?) shows a significantly higher proportion of ”Yes” responses in the selected agents. It suggests that LLM-based student simulations better capture students with such difficulties, aligning with real-world academic advising practices where these challenges are common. Notably, in the four trait questionnaire, the proportion of agents with strong self-efficacy and self-regulation increases substantially after filtering. It is because agents inherently display high self-efficacy and structured planning abilities in conversations, even though these traits may not typically align with students facing learning difficulties. Nevertheless, such agents still achieve high scores in behavior consistency evaluations.
@@ 4.5 Influence of Features in Profile
To investigate the influence of the features in the agent profile on the final profile and behavior scores, we performed a hot encoding of the features of the profile and used a Random Forest model [breiman2001random] to fit the propagated profile score and behavior score, where the size of test sets is 20% and the number of trees is 100. We then ranked the features based on their relative importance[archer2008empirical], where the relative importance of each feature is calculated by dividing its importance score by the highest among all.
@@ 4.6 Case Study
We showcase LLM scoring examples in Figure 8, comparing a good case and a bad case in profile and behavior evaluation, accompanied by expert scores, agreement levels, and explanations. In the good case (top), the questioning agent $Q$ detects an inconsistency in the student’s profile—17 years old while enrolled as a third-year Master’s student. The student agent fails to provide a compelling justification for this contradiction. Additionally, the LLM accurately assesses the student agent’s self-reported academic pressure, where the reasoning behind the discrepancy between daily and long-term pressure is somewhat reasonable. Regarding behavior evaluation, the scorer correctly identifies that the student agent violates the role of being a student who need academic advising. It demonstrates the model’s sensitivity to instruction following and reveals its occasional struggle with adherence to long instructions.
In contrast, the bad case (bottom) illustrates a scenario where the LLM’s evaluation deviates significantly from human judgment. The agent’s profile presents conflicting traits, such as experiencing stress while reporting no difficulty in tests and assignments. The LLM questions these inconsistencies but assigns a relatively low score due to its strict conflict detection. However, expert annotations suggest that such contradictions are not inherently erroneous, but reflect common real-world cases where students manage stress effectively despite academic challenges. Furthermore, the LLM scorer underperforms behavior evaluation by misinterpreting the agent’s conscientiousness and self-efficacy, leading to lower agreement scores with human experts. The results indicate that while LLM scorers is adept at identifying explicit conflicts, it may over-penalize cases where nuanced human explanations exist.
@@ 5 Conclusion
Our work explores the potential of automatically generating and filtering high-quality simulated student agents to enhance the authenticity of learning difficulty simulations. However, researchers and practitioners must consider several limitations and risks. First, our approach focuses on prompt-based student simulation, particularly extending the scope to metacognition. We do not explore more complex agent simulation frameworks like workflow-based or retrieval-augmented generation (RAG) simulations. These approaches will be explored in future work to improve simulation fidelity further. Second, some studies evaluate LLM-based human simulation by Turing tests [aher2023using], measuring the model’s success rate in mimicking human-like responses. Such evaluations emphasize linguistic style. Our work prioritizes metacognitive abilities and learning limitations in simulation rather than stylistic authenticity. We explicitly addressed this distinction during expert annotation to ensure the evaluation remains aligned with our research objectives. Third, our approach assumes that structured profiles can effectively capture student attributes and learning difficulties. However, real-world learning behaviors are highly dynamic and context-dependent. Future work should explore adaptive modeling techniques that better reflect the evolving nature of student cognition and engagement.
6 Acknowledgements
This work was supported by the SMP-Zhipu.AI Large Model Cross-Disciplinary Fund (Grant No. 20240211) and the National Natural Science Foundation of China (Grant No. 62407027).
7 Author Contributions
Haoxuan Li was responsible for the writing and revision of this paper, experimental design, and implementation; Jifan Yu, Xin Cong were responsible for advising the experimental design and writing the paper; Yang Dang was responsible for the experimental design; Daniel Zhang-Li, Lu Mi, Yisi Zhan, Huiqin Lin, Zhiyuan Liu were responsible for advising the writing and providing guidance on the research route of the paper. All authors approved the version to be published and agree to be accountable for all aspects of the work in ensuring that questions related to the accuracy or integrity of any part of the work are appropriately investigated and resolved.
Appendix A Ethical Considerations
@@
Simulated student agents are generated based on training data, which may introduce biases, stereotypes, or unintended inaccuracies in student representations. These biases could affect the fairness and applicability of educational interventions. Researchers must remain vigilant in mitigating such risks through continuous evaluation, bias detection, and responsible deployment of simulated agents.
@@ Appendix B Experimental Settings
B.1 Pipeline Performance Metrics
@@
$\text{NDCG@K}=\frac{1}{\text{IDCG@K}}\sum_{i=1}^{K}\frac{2^{\text{rel}(r_{i})}-1}{\log_{2}(i+1)},$ (8)